# Stochastic Optimal Control of an Industrial Power-to-Heat System with High-Temperature Heat Pump and Thermal Energy Storage

### v2

@@ 4.3 Optimal Quantizer for the Expected Value
The choice of the quantizer $\pazocal{Z}$ is important in order to obtain a good approximation of the expected value in (50). An approach was proposed by Pagès PagèsPrintems ; Pagès , in which a so-called optimal quantizer $\pazocal{Z}^{*}=\{z_{1}^{*},\ldots,z_{L}^{*}\}$ is selected. In the following, we will briefly discuss the optimality of the quantizer as well as some theoretical definitions and results. For more information, we refer the reader to the work of Pagès PagèsPrintems ; Pagès .
@@
Let $Z$ be a square integrable random variable in $\mathbb{R}^{2}$ and denote the probabilty density of $Z$ by $f_{Z}$. For the set $\pazocal{Z}=\{z_{1},\ldots,z_{L}\}\subset\mathbb{R}^{d}$, let $q:\mathbb{R}^{2}\to\pazocal{Z}$ be a Borel-measurable function. The random vector $q(Z)$ is called a $L$-quantization of $Z$ and $\pazocal{Z}$ is referred to as $L$-quantizer. The aim is to find a $L$-quantization $q$ such that the quadratic distortion $D_{L}^{Z}$ given by
@@
$q_{\text{Vor}}(z)=\sum_{l=1}^{L}z_{l}\mathbbm{1}_{C(z_{l})}(z)$ (57)
is minimizing $D_{L}^{Z}$, where $C(z_{l}),l=1,\ldots,L$ are Voronoi-cells with
$C(z_{l})\subset\{y\in\mathbb{R}^{d}|~{}\parallel z_{l}-y\parallel_{2}\leq\parallel z_{i}-y\parallel_{2},~{}i=1,\ldots,L\}.$ (58)
Note that the Voronoi-cells form a partition of $\mathbb{R}^{2}$. These quantizations can be understood as the nearest neighbor projection of $Z$ onto the set $\pazocal{Z}$. We denote the Voronoi $L$-quantization of $Z$ by $\widehat{Z}=q_{\text{Vor}}(Z)$. Moreover, the probability that $\widehat{Z}$ takes the value $z_{l}$ is given by
$p_{l}=\mathbb{P}(\widehat{Z}=z_{l})=\mathbb{P}(Z\in C(z_{l}))=\int\limits_{C(z_{l})}f_{Z}(z)\,\mathrm{d}z.$ (59)
@@
$D_{L}^{Z}(\pazocal{Z})=\mathbb{E}(\parallel Z-\widehat{Z}\parallel_{2}^{2})=\sum_{l=1}^{L}\mathbb{E}(\mathbbm{1}_{C(z_{l})}(Z)\parallel Z-z_{l}\parallel_{2}^{2})=\int\limits_{\mathbb{R}^{2}}f_{Z}(z)\underset{1\leq l\leq L}{\min\,}\parallel z-z_{l}\parallel_{2}^{2}\,\,\mathrm{d}z.$ (60)
Now for $\widehat{Z}$ the mapping $\pazocal{Z}\mapsto D_{L}^{X}(\pazocal{Z})$ is continuous and yields a minimum $\pazocal{Z}^{*}=\{z_{1}^{*},\ldots,z_{L}^{*}\}$ having distinct components PagèsPrintems . This set is called an optimal $L$-quantizer that satisfies
@@
Applying the optimal quantization to the Bellman equation (44), it is obvious to replace the continuous random variable $Z$ by its optimal quantization $\widehat{Z}$, associated with the optimal quantizer $\pazocal{Z}^{*}$ in order to obtain a reasonable approximation of the expected value. However, $g(\widehat{Z})$ is in general not an optimal quantizer for $g(Z)$, when applying the nonlinear transformation $g(z)=V_{n+1}(\pazocal{T}_{n}(x,a,z))$. If $g$ is bounded and continuous, then the convergence of $\mathbb{E}[\widehat{Z}]$ to $\mathbb{E}[Z]$ as the number of quantization points $L$ grows to infinity, see Pages2018 , and implies the convergence $\mathbb{E}[g(\widehat{Z})]\to\mathbb{E}[g(Z)]$. If we make further assumptions on $g$, the convergence in the sense
$\underset{L\to\infty}{\lim}L^{\alpha_{g}}|\mathbb{E}[g(Z)]-\mathbb{E}[g(\widehat{Z})]|\leq C_{g,Z},$
can be proven for different classes of functions $g$ and precise convergence rates $\alpha_{g}>0$ can be formulated, for more details see Lemaire2020 . The constant $C_{g,Z}$ depends on the properties of $g$ and the disturbance $Z$. In particular, if $g$ is a Lipschitz continuous function with Lipschitz coefficient $C_{L}$, we obtain that
$|\mathbb{E}[g(Z)]-\mathbb{E}[g(\widehat{Z})]|\leq C_{L}\mathbb{E}[\parallel Z-\widehat{Z}\parallel_{1}]\leq C_{L}\mathbb{E}[\parallel Z-\widehat{Z}\parallel_{2}]=C_{L}\sqrt{D^{Z}_{L}(\pazocal{Z}^{*})}.$ (62)
@@ 5 Reinforcement Learning Techniques
This section presents reinforcement learning algorithms that can tackle some of the problems mentioned in the context of BDP. We first introduce a quite general class of algorithms called temporal difference (TD) learning methods and then study Q-learning as a special case. Their aim is to approximate the value function in an appropriate parameter space and to construct the optimal policy with respect to this approximation. These methods rely on gradient descent to update the parameters with information obtained by generating samples of the controlled state process. In practice, the information for these methods does not have to come from an explicit model, as in our case. Instead, it can also be given as data from an observed real world process, which is why these algorithms are often referred to as model-free.
@@ 5.1 Temporal Difference Learning
In the following, we use a function approximation to approximate $V_{n}(x)$ for all $x\in\pazocal{X}$ and $n=0,\ldots,N-1$. Let $\theta_{n}\in\mathbb{R}^{p}$ be a parameter vector that describes an approximation of the exact value function $V_{n}$ in terms of $p\in\mathbb{N}$ parameters $\theta_{n}^{1},\cdots,\theta^{p}_{n}$
@@
Linear Function Approximation Konidaris_Osentoski_Thomas_2011 ; Sutton2018
In this class of functions, we represent $\overline{V}_{n}(x,\theta_{n})$ by a linear combination of basis or ansatz functions $\phi_{i},~{}i=1,\ldots,p$ with $\phi_{i}:\pazocal{X}\to\mathbb{R}$ as
@@
Feedforward neural networks (FNN) DeVore_Hanin_Petrova_2021 ; ELFWING20183
FNNs are a popular choice for nonlinear function approximators. Essentially, they consist of affine-linear maps and nonlinear activation functions. Let $d_{0}=\mathbb{R}^{d}$ and $d_{L}\in\mathbb{R}$ denote the input and output dimension of the FNN, then $\overline{V}_{n}(x,\theta_{n})$ is represented by the recursion
@@
where $L$ is the number of layers, $A_{l}\in\mathbb{R}^{d_{l}\times d_{l-1}}$ and $b_{l}\in\mathbb{R}^{d_{l}},~{}l=1,\ldots,L,$ are weights and biases for each layer with width $d_{l}\in\mathbb{N}$ and $\rho:\mathbb{R}\to\mathbb{R}$ is a nonlinear action function that is applied component-wise. The parameter vector $\theta_{n}$ is the collection of all matrices $A_{l}$ and vectors $b_{l}$. Examples for activation functions are the sigmoid function $\rho(x)=\frac{1}{1+e^{-x}}$ or the rectified linear unit (ReLU) $\rho(x)=\max\{x,0\}$. Nowadays, FNNs are frequently used because it is known that they fulfill the universal approximation property Hornik1989 , i.e. any continuous function can be approximated arbitrarily well.
@@
The underlying distribution in the expectation of loss (66) is called the state distribution and is used to sample states $X_{n}$. Normally, this distribution is chosen such that it reflects the importance of certain states that $X_{n}$ can take, i.e. states that are of interest for the controller or are likely to appear. A natural choice would be the distribution of $X_{n}$ or the so-called steady-state distribution Sutton2018 . For a given policy $u$, it describes the likelihood of $X_{n}$ taking a specific state for a given initial state. Sampling from the steady-state distribution is realized by creating trajectories starting from an initial state, while following the policy $u$. Here, the initial state is sampled from a predetermined distribution, for example a uniform distribution over the state space $\pazocal{X}$. Minimizing the loss with respect to $\theta_{n}$ can be achieved by gradient descent, which leads to an iterative update rule
@@
Here $k$ denotes the current iteration of the parameters and $\alpha_{n}^{k}>0$ is the step size or learning rate. Note that by interchanging the gradient with the expectation, we formally obtain
$\begin{split}\nabla_{\theta_{n}}\pazocal{L}(\theta_{n})&=\mathbb{E}[(V_{n}(X_{n})-\overline{V}_{n}(X_{n},\theta_{n}))\nabla_{\theta_{n}}\overline{V}_{n}(X_{n},\theta_{n})].\end{split}$ (68)
@@
with $\delta_{n}^{j}=V_{n}(x_{n}^{j})-\overline{V}_{n}(x_{n}^{j},\theta_{n}^{k})$. The iterative gradient update rule (69) is a special case of the Robbins-Monro algorithm Robbins1951 and is referred to as stochastic gradient descent (SGD) Garrigos2023 . By applying the Bellman equation (44), we have
@@
There are a several problems that need to be addressed before a parameter update can be performed.
@@
Firstly, since $V_{n+1}$ is unknown, we need to replace it with an approximation. One way to do this is to use a Monte-Carlo estimation of the performance criterion. Here, multiple trajectories starting from $X_{n}=x_{n}^{j}$ and thus multiple realizations of the performance criterion are obtained and averaged. However, the optimal policy is also unknown as it requires knowledge about the value function. Therefore, the best choice for the policy is the optimal policy induced by the value function approximation.
A more common approach is bootstrapping, which avoids following trajectories with a possibly suboptimal policy. In doing so, the value function $V_{n+1}$ is replaced by its corresponding parameterization $\overline{V}_{n+1}$. This brings some computational advantages, but at the expense of convergence properties toward the value function, which will be discussed below.
@@
Last but not least, as for BDP, there are several ways to calculate the expectation in (70). For BDP, a quantization approach is used, which could as well be used here. Nevertheless, a more common approach is to use a Monte-Carlo simulation to obtain sample of $X_{n+1}$ generated by the transition operator $\pazocal{T}_{n}$.
@@
$\begin{split}\theta_{n}^{k+1}=\theta_{n}^{k}-\alpha_{n}^{k}\frac{1}{M}\sum_{j=1}^{M}\delta_{n}^{j}\nabla_{\theta_{n}}\overline{V}_{n}(x_{n}^{j},\theta_{n}^{k})\end{split}$ (72)
@@ 5.2 Q-Learning
$\overline{Q}_{n}(x,a,\theta_{n}):\pazocal{X}\times\pazocal{A}_{n}(x)\times\mathbb{R}^{p}\to\mathbb{R}$ (74)
for approximating $Q^{*}_{n}(x,a)$. Due to the relation in (48), an approximate value function can also be derived by
@@
As for the TD-learning loss function (66), a state distribution for sampling states is used. Note that $u_{n}(X_{n})=A_{n}$ is a random variable that emphasizes the choice of an action distribution to obtain samples for $A_{n}$. For a given state $x$, this distribution can be directly defined by a probability measure on $\pazocal{A}_{n}(x)$ or by a selection policy $u^{S}=(u^{S}_{n})_{n=0,\ldots,N-1}$, with $A_{n}=u_{n}^{S}(X_{n})$. The relation (75) motivates to choose $u^{S}$ such that the optimal action is sampled frequently. A natural choice is a greedy selecting policy
@@
with respect to the current approximation. Analogous to TD-learning using (49) in loss (76) in combination with SGD, this now results in an iterative Q-learning update for the parameters $\theta_{n}$
@@
As above, we need to tackle similar problems to define a parameter update, such as replacing the unknown $Q_{n+1}^{*}$ and evaluating the expected value in (79). The most common variant of Q-learning uses bootstrapping to replace $Q_{n+1}^{*}$ and one-sample estimates for the expected value. This results in the temporal difference
@@
However, as a consequence of bootstrapping in TD-learning and Q-learning, the resulting gradient estimates may differ from those of the original underlying loss function (66). These kind of methods are different from SGD and are called as semi-gradient methods and require separate convergence results and additional assumptions. It should be noted that the main convergence analysis for TD-learning and Q-learning is based on MDPs with infinite time horizon. Convergence results for linear function approximation (64) are provided in Melo2008 , Melo2007 . Nonlinear function approximators such as neural networks require additional assumptions and techniques like projection NEURIPS2019_98baeb82 or linearization pmlr-v119-xu20c to guarantee convergence. Although these convergence results hold for infinite-horizon MDPs, they can still be applied to the finite-horizon case by augmenting the state with time as an additional state variable. Provided we can formulate an equivalent problem with the augmented state, the aforementioned convergence results can be applied.
@@
Given a state $X_{n}=x$, the selection policy $u^{S}$ is used to create samples of actions that ensure reasonable exploration of the action space $\pazocal{A}_{n}(x)$ and the state space of $X_{n+1}$. In the following, we discuss a commonly used class of selection policies that differ from the greedy policy (77). The greedy policy has one major disadvantage that can lead to poor approximations. If this greedy strategy is fully exploited, all decisions are based on the current function approximation, which itself is biased by approximation errors. This can lead to suboptimal action samples in the sense that the optimal action may not be sampled frequently. As a result, the approximation of $\overline{Q}$ will be poor as well as the value function approximation associated with it. We can compensate this by sampling actions from a uniform distribution on $\pazocal{A}_{n}(x)$. This approach has the advantage that all actions are selected with equal probability and an overall better approximation can be obtained for all actions. A drawback, however, is that the optimal action may again not get sampled very often. In the literature, this problem of choosing an appropriate $u^{S}$ is known as the exploration-exploitation dilemma. Here, we want to exploit the optimal policy induced by the function approximation as much as possible while still sampling a reasonable number of other actions to not miss out on the optimal action. In practice, a combination of random and greedy policies is used, known as $\varepsilon$-greedy policy with exploration rate $\varepsilon\in[0,1]$. Here, an action is either drawn from an uniform distribution on $\pazocal{A}_{n}(x_{n})$ with probability $\varepsilon$ or with probability $1-\varepsilon$ selected greedily as in (77). A simple linearly decaying exploration rate scheme is $\varepsilon^{k}=\frac{\varepsilon_{0}}{k}$ for $\varepsilon_{0}\in[0,1]$, where $k$ denotes the iteration counter.
@@
In Powell Powell2011 , examples of exploration rates and step sizes are discussed. It is also mentioned that choosing an appropriate step size that satisfies the Robbins-Monro conditions (81) is hard in practice. It may happen that the step sizes decrease too quickly, so the parameters converge to a non-optimal solution. Hence, it is suggested to use small constant step sizes $\alpha_{n}^{k}=\alpha_{0}>0$, as it is empirically observed that they work well in applications, although the second condition in (81) is violated.

### v3

@@ 4.3 Optimal Quantizer for the Expected Value
The choice of the quantizer $\pazocal{Z}$ is important in order to obtain a good approximation of the expected value in (51). An approach to obtain them was proposed by Pagès PagèsPrintems ; Pagès , in which a so-called optimal quantizer $\pazocal{Z}^{*}=\{z_{1}^{*},\ldots,z_{L}^{*}\}$ is selected. In the following, we will briefly discuss the optimality of this quantizer as well as some theoretical definitions and results. For more information, we refer the reader to the work of Pagès PagèsPrintems ; Pagès .
@@
For a square integrable random variable $Z$ in $\mathbb{R}^{2}$ with probability density of $Z$ by $f_{Z}$, we denote the set $\pazocal{Z}=\{z_{1},\ldots,z_{L}\}\subset\mathbb{R}^{2}$ and Borel-measurable function $q:\mathbb{R}^{2}\to\pazocal{Z}$. The random vector $q(Z)$ is called a $L$-quantization of $Z$ and $\pazocal{Z}$ is called $L$-quantizer. The aim is to find a $L$-quantization $q$ such that the quadratic distortion $D_{L}^{Z}$ given by
@@
$q_{\text{Vor}}(z)=\sum_{l=1}^{L}z_{l}\mathbbm{1}_{\pazocal{C}(z_{l})}(z)$ (57)
is minimizing $D_{L}^{Z}$, where $\pazocal{C}(z_{l}),l=1,\ldots,L$ are Voronoi-cells with
$\pazocal{C}(z_{l})\subset\{y\in\mathbb{R}^{d}|~\parallel z_{l}-y\parallel_{2}\leq\parallel z_{i}-y\parallel_{2},~i=1,\ldots,L\}.$ (58)
Note, that the Voronoi-cells form a partition of $\mathbb{R}^{2}$. These quantizations can be understood as the nearest neighbor projection of $Z$ onto the set $\pazocal{Z}$. We denote the Voronoi $L$-quantization of $Z$ by $\widehat{Z}=q_{\text{Vor}}(Z)$. Moreover, the probability that $\widehat{Z}$ takes the value $z_{l}$ is given by
$p_{l}=\mathbb{P}(\widehat{Z}=z_{l})=\mathbb{P}(Z\in\pazocal{C}(z_{l}))=\int\limits_{\pazocal{C}(z_{l})}f_{Z}(z)\,\mathrm{d}z.$ (59)
@@
$D_{L}^{Z}(\pazocal{Z})=\mathbb{E}(\parallel Z-\widehat{Z}\parallel_{2}^{2})=\sum_{l=1}^{L}\mathbb{E}(\mathbbm{1}_{\pazocal{C}(z_{l})}(Z)\parallel Z-z_{l}\parallel_{2}^{2})=\int\limits_{\mathbb{R}^{2}}f_{Z}(z)\underset{1\leq l\leq L}{\min\,}\parallel z-z_{l}\parallel_{2}^{2}\,\,\mathrm{d}z.$ (60)
Now for $\widehat{Z}$ the mapping $\pazocal{Z}\mapsto D_{L}^{X}(\pazocal{Z})$ is continuous and yields a minimum $\pazocal{Z}^{*}=\{z_{1}^{*},\ldots,z_{L}^{*}\}$ having distinct components PagèsPrintems . This set is called an optimal $L$-quantizer and satisfies
@@
When applying the optimal quantization to (44), it is obvious to replace the continuous random variable $Z$ by its optimal quantization $\widehat{Z}$, associated with the optimal quantizer $\pazocal{Z}^{*}$ in order to obtain a reasonable approximation of the expected value. However, g(Z^)g($\widehat${Z}) is in general not an optimal quantizer for $g(Z)$, when applying the nonlinear transformation $g(z)=V_{n+1}(\pazocal{T}_{n}(x,a,z))$. If $g$ is bounded and continuous, then the convergence of $\mathbb{E}[\widehat{Z}$] to $\mathbb{E}[Z]$ as the number of quantization points $L$ grows to infinity, see Pages2018 , implies the convergence $\mathbb{E}[g(\widehat{Z})]\to\mathbb{E}[g(Z)]$. If we make further assumptions on $g$, the convergence in the sense
$\underset{L\to\infty}{\lim}L^{\alpha_{g}}|\mathbb{E}[g(Z)]-\mathbb{E}[g(\widehat{Z})]|\leq C_{g,Z},$ (62)
can be proven for different classes of functions $g$ and precise convergence rates $\alpha_{g}>0$ can be formulated, for more details see Lemaire2020 . The constant $C_{g,Z}$ depends on the properties of $g$ and the disturbance $Z$. In particular, if $g$ is a Lipschitz continuous function with Lipschitz coefficient $L_{g}$, we obtain that
$|\mathbb{E}[g(Z)]-\mathbb{E}[g(\widehat{Z})]|\leq L_{g}\mathbb{E}[\parallel Z-\widehat{Z}\parallel_{1}]\leq L_{g}\mathbb{E}[\parallel Z-\widehat{Z}\parallel_{2}]=L_{g}\sqrt{D^{Z}_{L}(\pazocal{Z}^{*})}.$ (63)
@@ 5 Reinforcement Learning Techniques
This section presents reinforcement learning algorithms that can tackle some of the problems mentioned in the context of BDP. We first introduce a quite general class of algorithms called temporal difference (TD) learning methods and then study Q-learning as a special case. Their aim is to approximate the value function in an appropriate parameter space and to construct the optimal policy with respect to this approximation. These methods rely on gradient descent to update the parameters with information obtained by generating samples of the controlled state process. In practice, the information for these methods does not have to come from an explicit model, as in our case. Instead, it can also be provided as data from an observed real world process, which is why these algorithms are often referred to as model-free.
@@ 5.1 Temporal Difference Learning
In the following, we use a function approximation to approximate $V_{n}(x)$ for all $x\in\pazocal{X}$ and $n=0,\ldots,N-1$. Let $\theta_{n}\in\mathbb{R}^{p}$ be a parameter vector that describes an approximation $\overline{V}_{n}$ of the exact value function $V_{n}$ in terms of $p\in\mathbb{N}$ parameters $\theta_{n}^{1},\cdots,\theta^{p}_{n}$
@@
Linear Function Approximation
In this class of functions Konidaris_Osentoski_Thomas_2011 ; Sutton2018 , $\overline{V}_{n}(x,\theta_{n})$ is represented by a linear combination of basis or ansatz functions $\phi_{i},~i=1,\ldots,p$ with $\phi_{i}:\pazocal{X}\to\mathbb{R}$ as
@@
Feedforward Neural Networks
Feedforward neural networks (FNN) are simple artificial neural networks and are a popular choice for nonlinear function approximators DeVore_Hanin_Petrova_2021 ; ELFWING20183 . Essentially, they consist of affine-linear maps and nonlinear activation functions. Let $d_{0}=\mathbb{R}^{d}$ and $d_{L}\in\mathbb{R}$ denote the input and output dimension of the FNN, then $\overline{V}_{n}(x,\theta_{n})$ is represented by the recursion
@@
where $L$ is the number of layers, $A_{l}\in\mathbb{R}^{d_{l}\times d_{l-1}}$ and $b_{l}\in\mathbb{R}^{d_{l}},~l=1,\ldots,L,$ are weights and biases for each layer with width $d_{l}\in\mathbb{N}$ and $\rho:\mathbb{R}\to\mathbb{R}$ is a nonlinear activation function that is applied component-wise. The parameter vector $\theta_{n}$ is the collection of all matrices $A_{l}$ and vectors $b_{l}$. Examples for activation functions are the sigmoid function $\rho(x)=\frac{1}{1+e^{-x}}$ or the rectified linear unit (ReLU) $\rho(x)=\max\{x,0\}$. Nowadays, FNNs are frequently used because it is known that they fulfill the universal approximation property Hornik1989 , i.e., any continuous function can be approximated arbitrarily well.
@@
The underlying distribution in the expectation of loss (67) is called the state distribution and is used to sample states $X_{n}$. Normally, this distribution is chosen such that it reflects the importance of certain states that $X_{n}$ can take, i.e., states that are of interest for the controller or are likely to appear. A natural choice would be the distribution of $X_{n}$ or the so-called steady-state distribution Sutton2018 . For a given policy $u$, it describes the likelihood of $X_{n}$ taking a specific state for a given initial state. Sampling from the steady-state distribution is realized by creating trajectories starting from an initial state, while following the policy $u$. Here, the initial state is sampled from a predetermined distribution, for example a uniform distribution over the state space $\pazocal{X}$. Minimizing the loss with respect to $\theta_{n}$ can be achieved by gradient descent, which leads to an iterative update rule
@@
Here $k$ denotes the current iteration of the parameters and $\alpha_{n}^{k}>0$ is the step size or learning rate. Note, that by interchanging the gradient with the expectation, we formally obtain
$\begin{split}\nabla_{\theta_{n}}\pazocal{L}(\theta_{n})&=\overline{\mathbb{E}}[(V_{n}(X_{n})-\overline{V}_{n}(X_{n},\theta_{n}))\nabla_{\theta_{n}}\overline{V}_{n}(X_{n},\theta_{n})].\end{split}$ (69)
@@
with $\delta_{n}^{j}=V_{n}(x_{n}^{j})-\overline{V}_{n}(x_{n}^{j},\theta_{n}^{k})$. The iterative gradient update rule (70) is a special case of the Robbins-Monro algorithm Robbins1951 and is referred to as stochastic gradient descent (SGD) Garrigos2023 . By applying (44), we get
@@
There are some problems that need to be addressed before a parameter update can be performed.
@@
Firstly, since $V_{n+1}$ is unknown, we need to replace it with an approximation. One way to do this is to use a Monte-Carlo estimation of the performance criterion. Here, multiple trajectories starting from $X_{n}=x_{n}^{j}$ and thus multiple realizations of the performance criterion are obtained and averaged. However, the optimal policy is also unknown as it requires knowledge about the value function. Thus, the best choice for the policy is the (sub)optimal policy induced by the value function approximation.
A more common approach is bootstrapping, which avoids following trajectories with a possibly suboptimal policy. In doing so, the value function $V_{n+1}$ is replaced by its corresponding parameterization $\overline{V}_{n+1}$. This brings some computational advantages, but at the expense that convergence towards the value function can not always be shown, which will be discussed below.
@@
Last but not least, as for BDP, there are several ways to calculate the expectation in (71). For BDP, a quantization approach is used, which could as well be used here. Nevertheless, a more common approach is to use a Monte-Carlo simulation and use samples of $X_{n+1}$.
@@
$\begin{split}\theta_{n}^{k+1}=\theta_{n}^{k}-\alpha_{n}^{k}\frac{1}{M}\sum_{j=1}^{M}\delta_{n}^{j}\nabla_{\theta_{n}}\overline{V}_{n}(x_{n}^{j},\theta_{n}^{k}),\end{split}$ (73)
@@ 5.2 Q-Learning
$\overline{Q}_{n}(x,a,\theta_{n}):\pazocal{X}\times\pazocal{A}\times\mathbb{R}^{p}\to\mathbb{R},$ (75)
in order to approximate $Q^{*}_{n}(x,a)$. Due to the relation in (49), an approximate value function can also be derived by
@@
As for the TD-learning loss function (67), a sample state distribution is used. Note, that $u_{n}(X_{n})=A_{n}$ is a random variable, which emphasizes the choice of an action distribution to obtain samples for $A_{n}$. For a given state $x$, this distribution can be directly defined by a probability measure on $\pazocal{A}_{n}(x)$ or by a selection policy $u^{S}=(u^{S}_{n})_{n=0,\ldots,N-1}$, with $A_{n}=u_{n}^{S}(X_{n})$. The relation (76) motivates to choose $u^{S}$ such that the optimal action is sampled frequently. A natural choice is a greedy selecting policy
@@
with respect to the current approximation. Analogous to TD-learning using (50) in loss (77) in combination with SGD, this results in an iterative Q-learning update for the parameters $\theta_{n}$
@@
As above, we need to tackle similar problems to define a parameter update, such as replacing the unknown $Q_{n+1}^{*}$ and evaluating the expected value in (80). The most common variant of Q-learning uses bootstrapping to replace $Q_{n+1}^{*}$ and one-sample estimates for the expected value, which results in the temporal difference
@@
However, as a consequence of bootstrapping in TD-learning and Q-learning, the resulting gradient estimates may differ from those of the original underlying loss function (67). These kind of methods are different from SGD and are called semi-gradient methods and require separate convergence results as well as additional assumptions. It should be noted that the main convergence analysis for TD-learning and Q-learning is based on MDPs with infinite time horizon. Convergence results for linear function approximation (65) are provided in Melo2008 , Melo2007 . Nonlinear function approximators such as neural networks require additional assumptions and techniques like projection NEURIPS2019_98baeb82 or linearization pmlr-v119-xu20c to guarantee convergence. Although these convergence results hold for infinite-horizon MDPs, they can still be applied to the finite-horizon case by augmenting the state with time as an additional state variable. Provided that we can formulate an equivalent problem with the augmented state, the aforementioned convergence results can be applied.
@@
Given a state $X_{n}=x$, the selection policy $u^{S}$ is used to create samples of actions that ensure reasonable exploration of the action space $\pazocal{A}_{n}(x)$ and the state space of $X_{n+1}$. In the following, we discuss a commonly used class of selection policies that differ from the greedy policy (78). The greedy policy has one major disadvantage that can lead to poor approximations. If this greedy strategy is fully exploited, all decisions are based on the current function approximation, which itself is biased by approximation errors. This can lead to suboptimal action samples in the sense that the optimal action may not be sampled frequently. As a result, the approximation of $\overline{Q}$ will be poor as well as the value function approximation associated with it. We can compensate this by sampling actions from a uniform distribution on $\pazocal{A}_{n}(x)$. This approach has the advantage that all actions are selected with equal probability and an overall better approximation can be obtained for all actions. A drawback, however, is that the optimal action may again not get sampled very often. In the literature, this problem of choosing an appropriate $u^{S}$ is known as the exploration-exploitation dilemma. Here, we want to exploit the optimal policy induced by the function approximation as much as possible while still sampling a reasonable number of other actions to not miss out on more optimal actions. In practice, a combination of random and greedy policies is used, known as $\varepsilon$-greedy policy with exploration rate $\varepsilon\in[0,1]$. Here, an action is either drawn from an uniform distribution on $\pazocal{A}_{n}(x_{n})$ with probability $\varepsilon$ or with probability $1-\varepsilon$ selected greedily as in (78). The exploration rate could be given by a simple linearly decaying scheme with $\varepsilon^{k}=\frac{\varepsilon_{0}}{k}$ for $\varepsilon_{0}\in[0,1]$, where $k$ denotes the iteration counter.
@@
In Powell Powell2011 , examples of exploration rates and step sizes are discussed. It is also mentioned that choosing an appropriate step size that satisfies the Robbins-Monro conditions (82) is hard in practice. It may happen that the step sizes decrease too quickly, so the parameters converge to a non-optimal solution. Hence, it is suggested to use small constant step sizes $\alpha_{n}^{k}=\alpha_{0}>0$ as it has been empirically observed that these work well in applications, although the second condition in (82) is violated.